# Humanoid Locomotion and Manipulation: Current Progress and Challenges in Control, Planning, and Learning thanks: *co-corresponding authors

### v1

@@ III Tactile Sensing
Humanoid locomotion and manipulation involve extensive physical interactions with the environment and objects, requiring multimodal sensing for understanding the environment, tracking manipulated objects, and evaluating how contact impacts the balance of both the robot and the objects. Visual sensors have shown effectiveness in object tracking and simultaneous localization and mapping (SLAM) [84], while proprioceptive sensors are usually combined to estimate contact information in contact-rich tasks [80]. These sensory modalities have been widely adopted in existing systems and have been thoroughly reviewed in the literature [12]. This survey complements existing research by focusing on a less explored but equally critical sensing modality: tactile sensing.
@@
Numerous studies have developed tactile sensors based on various transduction principles that can sense normal and tangential forces, vibration, temperature, and pre-contact proximity information. Comparative studies of various sensor designs can be found in [90, 91, 92]. This survey instead focuses on their application in humanoid loco-manipulation, categorized into three areas: (i) tactile sensing on hands, (ii) tactile sensing on foot soles, and (iii) whole-body tactile sensing. The following sections review recent advancements in each domain, emphasizing their roles in balancing control, scheduling contacts, and enhancing interaction capabilities, as illustrated in Fig. 4.
@@ III-A Tactile Sensing on Hands
More complex in-hand manipulation tasks demand interactive perception beyond static object models. Dynamic contact information, including real-time tracking of object states, monitoring contact stability [105], and predicting interaction outcomes [94], i.e., how contact forces affect the balance of both objects and the robot, are crucial to achieving complex interactive behaviors. However, due to the inherent complexity of multi-contact dynamics and increased dimensionality of the contact state space, model-based methods still struggle to match human-level dexterity and versatility in multi-finger manipulation.
Alternatively, model-free Reinforcement Learning (RL) has shown the ability to address complex contact interactions. These approaches integrate tactile measurements directly into the state space to train end-to-end policies [106, 93]. Tactile-based RL faces two main challenges: (i) the high-dimensional input space of raw tactile sensor data and (ii) the difficulty of accurately simulating contact physics for sim-to-real transfer. To tackle these challenges, dimensionality reduction techniques such as spectral clustering, principal component analysis [107] and autoencoders [108] have been explored, while [93] employs Deep Reinforcement Learning (DRL) to manage high-dimensional input space. Moreover, there have been increasing efforts in developing tactile simulators [109, 110] to improve the accessibility of simulated tactile data and facilitate zero-shot sim-to-real transfer [111].
Besides task-specific RL, other learning methods have been sought for more generalizable policies. [112] employs diffusion policy to achieve complex and long-horizon bi-manual manipulation tasks, while recent work has integrated tactile sensing into foundation models alongside vision and language [113, 114]. Though limited to simple control tasks, these models may eventually enable more natural and versatile physical interactions in humanoid robots.
@@ III-B Tactile Sensing on Feet
Besides manipulation, tactile sensing has started to gain traction for locomotion problems. For legged locomotion, estimation of Ground Reaction Forces (GRFs) and terrain properties is critical for maintaining whole-body stability on diverse, uneven surfaces. While vision and proprioception sensors can provide an indirect estimation of the terrain, these sensing modules lack the capability of accurately estimating GRFs and various terrain properties. Tactile sensing on foot soles has the potential to provide direct, unobstructed, and accurate contact measurements, but remains largely underexplored.
@@
To date, tactile sensors for legged systems have been mainly applied for monopods, quadrupeds, and hexapods, enabling the functionalities of classifying terrain [121, 122, 123], detecting contact forces and soil flow [124], detecting contact angle [125] and type (e.g., surface, edge, or no contact) [126, 127], and estimating 3D pressure distribution [123].
@@
Few studies have built tactile sensors for humanoid robots. These sensors are mostly used for applications including terrain classification [128, 97] and ground slope recognition [118]. The sensed tactile information should aid the control of humanoid dynamics and enhance the locomotion performance. A notable work in this direction [96] reconstructs the pressure shape of the foothold, enabling the recognition of uneven terrain and replanning footsteps in real time.
@@ III-C Whole-body Tactile Sensing
Large-area tactile sensing significantly enhances a robot’s ability to handle large objects, including object identification through tactile exploration and whole-body manipulation. For example, [130] enable a humanoid robot NAO, covered with artificial skin over its entire upper body, to classify large, heavy objects with different weights and textures. [88] demonstrate whole-arm tactile sensing by reaching objects in cluttered spaces while regulating contact forces across its arms. Close-proximity whole-body capacitive sensing is implemented in [131], enabling a cobotic humanoid with workers close-presence awareness. The same technology is used to draw semantics in human-humanoid physical interaction in [132]. Moving away from traditional methods [133] that prioritize collision-avoiding trajectories, [134] utilize tactile feedback to detect and clear movable obstacles, thereby solving the problem of navigation among movable objects. Additionally, with tactile sensors covering their arms and chest, humanoid robots HRP-2 and Punyo-1 can use their entire upper body to grasp and lift large, heavy boxes [135, 136] or various household items [62].
@@
However, current works of tactile-based whole-body manipulation are still limited to grasping or simple pick-and-place motions with the upper body. This is due to the significant challenges of whole-body manipulation, as discussed in Sec. II-C, including understanding the complex contact dynamics of multiple contacts, handling the high dimensionality of the sensor data, and addressing the sim-to-real gap.
@@
Despite the promising potential that tactile sensors provide, human-like loco-manipulation with more dynamic interactions and contact shifting, such as transferring weight to one arm to free the other for tasks like opening a door, requires sophisticated integration of all system levels in perception, planning, and control. A major challenge in tactile perception, and a hurdle for a tight integration with planning and control, is the difficulty in dynamically reasoning about contacts. This involves not only estimating contact points and static object models but, more crucially, understanding how these contacts and changes of contacts impact the system in real-time, including the balance of both the robot and the object. Such information is vital for a planner to make informed decisions and, in a learning framework, can enhance sample efficiency.
@@
Conclusion: Tactile sensing is yet an underexplored modality for advancing humanoid loco-manipulation, providing direct contact information necessary for tasks involving complex interactions with environments and objects. While tactile perceptions have significantly enhanced humanoid tasks, achieving human-level dexterity and versatility remains challenging, requiring further research into dynamic perception, and multimodal sensing integration to enable systematic, real-time decision-making during interactions. This includes optimizing contact scheduling on whole-body based on object properties like size and weight, and understanding how contact dynamics affect robot and object balance during simultaneous loco-manipulation. Moreover, the design of whole-body tactile systems should account for varying sensor resolutions and load requirements, i.e. hands need higher resolution for delicate tasks, while body skin can operate at lower resolution but withstand higher payloads. For further reading, we recommend a survey paper on humanoid tactile sensing [90], and a book chapter on tactile sensing technologies with an emphasis on deployment on humanoid robot [137].
@@ IV Multi-Contact Planning for Humanoids
Multi-contact planning remains one of the most challenging tasks in robotics. Specifically, in the context of humanoid whole-body loco-manipulation, a planner ought to solve trajectories that handle rich interactions with environments or objects. Particularly, besides robot state trajectories, the planner is also expected to decide contact position (or contact location), contact mode, and contact force in a loco-manipulation task. Given the underactuated nature of humanoid robots and the addition of manipulation interaction dynamics, maintaining balance and manipulating objects rely solely on these contact interactions, which already makes multi-contact planning a challenging problem. Moreover, the diverse physical properties of environments and objects (e.g., rigid or soft, fixed or movable) complicate the problem even further.
Over the past decade, the field has produced fruitful results in multi-contact humanoid planning, demonstrating promising potential across various locomotion and manipulation tasks [138, 139, 3, 140]. However, these works require pre-planned contact mode sequences before planning robot whole-body motion trajectories [141], leaving an open problem: how to solve the locomotion and manipulation contact planning problem simultaneously with the whole-body trajectory planning in a unified fashion, a.k.a, Contact-Implicit Planning (CIP) [142, 143]. The primary challenge of this CIP lies in its high computational burden and combinatorial complexity of identifying potential contact mode sequences. Therefore, selecting suitable approaches depends on the specific problem requirements, including factors such as solving time, numerical robustness of the solution, resolution of the solution, and dependency on numerical models.
To choose appropriate multi-contact planning algorithms for a high degrees-of-freedom underactuated system, state-of-the-art approaches present three main categories: (i) searching, (ii) optimization, and (iii) learning, as illustrated in Fig. 5.
@@ IV-A Search-based Contact Planning
Search-based approaches employ state expansion that allows exploring configurations to make and break contacts; collisions and kinematic feasibility are often checked during each search step. Heuristics can be applied in search-based methods for efficient exploration. The search result is an optimal sequence of contact modes that ensure stability and task efficiency. Whole-body motions can be optimized during the search to verify the dynamic feasibility of candidate contact sequences [144] or after the search in a contact-before-motion style [145]. Search-based methods are commonly used for gait planning in legged robot locomotion [146, 147, 148, 149]. Expanding their capabilities in more intricate multi-contact loco-manipulation planning, [150] implement a graph search method for humanoid grasp contact planning and replanning. [151] introduce a contact-before-motion planner for multi-contact behaviors.
@@ IV-B Optimization-based Contact Planning
Due to the inherently large problem size, CITOs often rely on speed-up strategies, such as warm-starting with reasonable initial guesses for fast convergence [162] and separating into contact planning and whole-body motion planning subproblems in a hierarchical fashion [163]. With the increasing demand for computation efficiency, CITOs have witnessed a rise in computation speed via sequential quadratic programming (e.g.[142]), differential dynamic programming (e.g.[164]), and iterative linear quadratic regulator (e.g.[165]). These improvements have even enabled the use of CITO in a Model Predictive Control (a.k.a. CI-MPC) framework for real-time planning on quadruped robots [166, 167] and robotic arms [168, 169]. However, for the humanoid robots, applying CITO to loco-manipulation has yet to be achieved.
Migrating such CITO as real-time CI-MPC to humanoid loco-manipulation problems presents its own set of challenges, including high-dimensional space of optimization variables, complex/undifferentiable contact dynamics models, proper modeling of interaction dynamics, resolution of initial guess, and tedious tuning. Although the body of literature on CITO and CI-MPC for humanoid robots is quite limited, recent efforts have demonstrated viable strategies for accelerating such methods. To achieve real-time CITO in multi-contact humanoid motion generation, [164] improves the time stepping integrator-based method [170] by introducing a smooth-max function to approximate the contact impulses. [171] incorporates dynamic complementarity conditions in the rigid contact model and improved solving time in contact implicit humanoid locomotion problems. [166] leverages structure exploitation and offsets the time-consuming Linear Complementarity Problems (LCPs) by pre-computing its constant terms offline for improved online computation efficiency.
Furthermore, the potential to harness and combine the advantages of search-based and optimization-based methods remains largely unexplored. For example, [172] enhanced CITO by incorporating a graph-search-based contact sequence generator and neural-network-based capturability prediction for efficient and robust disturbance rejection in humanoid multi-contact locomotion. To improve the robustness of CITO, [173, 174] also show that considering the uncertainty in the optimization results in solutions robust against perturbation from the terrain contact.
@@ IV-C Learning-based Contact Planning
In addition to search-based and optimization-based methods, learning-based approaches have demonstrated promising potential in planning for multi-contact tasks, such as using reinforcement learning to plan for velocity commands and contact sequences [175, 176]. These learning-based planners are mostly modular, making it possible to form a hierarchical architecture with model-based planners and controllers at the low level. Compared with traditional optimization-based or heuristics-based approaches, learning-based elements enhance the computation efficiency in multi-contact planning. For example, [177] learn a prediction of centroidal dynamics evolution for efficient contact sequence generation under $0.1$ s, allowing a $300$ times computation speed boost compared to traditional optimization-based methods.
@@
In addition, learning-based approaches can assist contact prediction, which allows additional information for contact (re)planning in real time. Precise contacts are often hard to obtain from motion capture data, making it challenging to learn directly from data. To synthesize plausible motion, a naive supervised learning approach often leads to objects moving without any contact or significant penetration between the predicted human body and the objects. [178] introduced contact correction and predicts motions relative to the contacts predicted. [179] separate the contact prediction and whole-body motion prediction by first predicting the contact positions of a moving object, which are then used as a constraint to synthesize whole-body motion. These models have the potential to serve as a loco-manipulation planner for humanoid robots. [180] learn to find from the video scenes the affordance (i.e., potential contact points for task execution). These contacts can be used as heuristics for subsequent motion planning.

### v2

@@ III Tactile Sensing
Humanoid locomotion and manipulation involve extensive physical interactions with the environment and objects, requiring multimodal sensing for understanding the environment, tracking manipulated objects, and evaluating how contact impacts the balance of both the robot and the objects. Visual sensors have shown effectiveness in object tracking and simultaneous localization and mapping (SLAM) [93], while proprioceptive sensors are usually combined to estimate contact information in contact-rich tasks [89]. These sensory modalities have been widely adopted in existing systems and have been thoroughly reviewed in the literature [19].
In contrast, tactile sensing, despite being uniquely suited for capturing detailed contact information critical for both locomotion and manipulation, has received relatively limited attention in existing humanoid literature. This survey specifically highlights tactile sensing to address this gap and to underscore its complementary role alongside other sensing modalities in humanoid systems.
@@
Numerous studies have developed tactile sensors based on various transduction principles that can sense normal and tangential forces, vibration, temperature, and pre-contact proximity information. Comparative studies of various sensor designs can be found in [13, 99, 100]. This survey instead focuses on their application in humanoid loco-manipulation, categorized into three areas: (i) tactile sensing on hands, (ii) tactile sensing on foot soles, and (iii) whole-body tactile sensing. The following sections review recent advancements in each domain, emphasizing their roles in balancing control, scheduling contacts, and enhancing interaction capabilities, as illustrated in Fig. 4 and Table II.
TABLE II: Roles and challenges of tactile sensing in locomotion and manipulation applications
Application Roles of Tactile Sensing Body Region Involved Challenges Solutions
Manipulation
• Grasping force and stability regulation
[101, 102, 103, 104]
• Object property and state estimation
[105, 106, 107]
• Contact dynamics estimation
[108, 109]
• Input modality for RL/IL/FM training
[110, 111, 112, 113, 114]
• Hand (in IHM)
• Whole-body (in WHM)
• High-dimensional
contact space
• Tactile physics simulation
• Sim-to-real transfer
• Dimensionality reduction
[115, 116, 111]
• Tactile simulators
[117, 118, 119]
Locomotion
• GRF and support region estimation
[120]
• Terrain property recognition
[121, 122]
• Feet
• Sensor durability against
heavy weight and various
terrains
• Durable materials
and mechanical design
[120, 121, 122]
Loco-manipulation
• All of above
• Unified locomotion and manipulation
planning
• Whole-body
• All of Above
• Real-time contact
reasoning and adaptation
• To be explored
IHM - In-hand Manipulation;   WBM - Whole-body Manipulation.
@@ III-A Tactile Sensing on Hands
More complex in-hand manipulation tasks demand interactive perception beyond static object models. Dynamic contact information, including real-time tracking of object states, monitoring contact stability [106], and predicting interaction outcomes [108, 109], i.e., how contact forces affect the balance of both objects and the robot, are crucial to achieving complex interactive behaviors. However, due to the inherent complexity of multi-contact dynamics and increased dimensionality of the contact state space, model-based methods still struggle to match human-level dexterity and versatility in multi-finger manipulation.
Alternatively, model-free Reinforcement Learning (RL) has shown the ability to address complex contact interactions. These approaches integrate tactile measurements directly into the state space to train end-to-end policies [110, 111]. Other learning methods besides task-specific RL have been sought for more generalizable policies. [112] employs diffusion policy to achieve complex and long-horizon bi-manual manipulation tasks, while recent work has integrated tactile sensing into foundation models alongside vision and language [113, 114]. Though limited to simple control tasks, these models may eventually enable more natural and versatile physical interactions in humanoid robots.
@@ III-B Tactile Sensing on Feet
Other than manipulation, tactile sensing has started to gain traction for locomotion problems. For legged locomotion, estimation of Ground Reaction Forces (GRFs) and terrain properties is critical for maintaining whole-body stability on diverse, uneven surfaces. While vision and proprioception sensors can provide an indirect estimation of the terrain, these sensing modules lack the capability of accurately estimating GRFs and various terrain properties. Tactile sensing on foot soles has the potential to provide direct, unobstructed, and accurate contact measurements, but remains largely underexplored.
@@
To date, tactile sensors for legged systems have been primarily developed for monopods, quadrupeds, and hexapods. Only a few studies have built tactile feet sensors for humanoid robots, with applications such as terrain classification [121] and ground slope recognition [122]. The sensed tactile information should aid the humanoid control and enhance the locomotion performance. A notable work in this direction [120] reconstructs the pressure shape of the foothold, enabling the recognition of uneven terrain and footstep replanning in real time.
@@ III-C Whole-body Tactile Sensing
Large-area tactile sensing significantly enhances a robot’s ability to handle large objects, including object identification through tactile exploration and whole-body manipulation. For example, [107] enables a humanoid robot NAO, covered with artificial skin over its entire upper body, to classify large, heavy objects with different weights and textures. [97] demonstrates whole-arm tactile sensing by reaching objects in cluttered spaces while regulating contact forces across its arms. Close-proximity whole-body capacitive sensing is implemented in [134], enabling a cobotic humanoid with workers close-presence awareness. The same technology is used to draw semantics in human-humanoid physical interaction in [135]. Moving away from traditional methods [136] that prioritize collision-avoiding trajectories, [137] uses tactile feedback to detect and clear movable obstacles, thereby solving the problem of navigation among movable objects. Additionally, with tactile sensors covering their arms and chest, humanoid robots HRP-2 and Punyo-1 can use their entire upper body to grasp and lift large, heavy boxes [138, 104] or various household items [71].
III-D Challenges in Tactile Sensing
@@
Current works of tactile-based manipulation, whether in-hand or whole-body, are still largely limited to grasping or simple motions like rolling, pushing, or pick-and-place. Although tactile sensing holds great promise, its integration into more dexterous manipulation or loco-manipulation tasks that involve more dynamic interactions and contact shifting over the full body continues to pose substantial challenges. These include understanding the complex multi-contact dynamics, handling the high dimensionality of the sensor data, and addressing the sim-to-real gap.
Several approaches have been explored to address these difficulties. For instance, dimensionality reduction techniques such as spectral clustering, principal component analysis [115], and autoencoders [116] have been employed to manage the high-dimensional input space. Others have proposed simplifying sensor outputs by using boolean contact states (i.e., contact or no-contact) instead of continuous force values in RL training [111]. Furthermore, advances in tactile simulation [117, 118] are making simulated tactile data more accessible, thereby supporting zero-shot sim-to-real transfer [119].
@@
One of the remaining fundamental challenges is to dynamically reason about contacts. This involves not only estimating contact states and static object models but, more crucially, understanding how these contacts and changes of contacts impact the system in real time, including the balance of both the robot and the object. Such information is vital for a planner to make informed decisions and, in a learning framework, can enhance sample efficiency. Addressing this challenge calls for deeper collaboration between the sensing and control communities to determine what tactile information should be abstracted and how to tightly integrate real-time tactile perception into control and planning for humanoid tasks.
@@
Key Takeaways: Tactile sensing is yet an underexplored modality for advancing humanoid loco-manipulation, providing direct contact information necessary for tasks involving complex interactions with environments and objects. While it has improved task performance, achieving human-level dexterity and versatility requires advances in dynamic perception, and multi-modal sensing integration to enable systematic, real-time decision-making during interactions. Future directions include optimizing whole-body contact scheduling based on object properties and understanding how contact dynamics affect robot and object balance during loco-manipulation. Moreover, the design of whole-body tactile systems should account for varying sensor resolutions and load requirements, i.e. hands need higher resolution for delicate tasks, while body skin can operate at lower resolution but withstand higher payloads. For further reading, we recommend a survey paper on humanoid tactile sensing [13], and a book chapter on tactile sensing technologies with an emphasis on deployment on humanoid robot [139].
@@ IV Multi-Contact Planning for Humanoids
Multi-contact planning remains one of the most challenging tasks in robotics. Specifically, in the context of humanoid whole-body loco-manipulation, a planner ought to solve trajectories that handle rich and intricate interactions with environments or objects. Particularly, beyond robot state trajectories, the planner is also expected to decide contact position (or contact location), contact mode, and contact force in a loco-manipulation task. Given the underactuated nature of humanoid robots and the addition of manipulation interaction dynamics, maintaining balance and manipulating objects rely solely on these contact interactions. This already makes multi-contact planning a challenging problem. Moreover, the diverse physical properties of environments and objects (e.g., rigid or soft, fixed or movable) complicate the problem even further.
Over the past decade, the field has produced fruitful results in multi-contact humanoid planning, demonstrating promising potential across various locomotion and manipulation tasks [140, 141, 3, 142]. However, these works require pre-planned contact mode sequences before planning robot whole-body motion trajectories [143]. This raises an open problem: how to solve the locomotion and manipulation contact planning problem simultaneously with the whole-body trajectory planning in a unified fashion, a.k.a, Contact-Implicit Planning (CIP) [144, 145]. The primary challenge of this CIP lies in its high computational burden and combinatorial complexity of identifying potential contact mode sequences. Therefore, selecting suitable approaches depends on the specific problem requirements, including factors such as solving time, numerical robustness of the solution, resolution of the solution, and dependency on numerical models.
State-of-the-art procedures select planning algorithms for the underactuated system with three main categories: (i) searching, (ii) optimization, and (iii) learning, as illustrated in Fig. 5.
@@ IV-A Search-based Contact Planning
Search-based approaches employ state expansion that allows exploring configurations to make and break contacts; collisions and kinematic feasibility are often checked during each search step. Heuristics can be applied in search-based methods for efficient exploration. The search result is an optimal sequence of contact modes that ensure stability and task efficiency. Whole-body motions can be optimized during the search to verify the dynamic feasibility of candidate contact sequences [146] or after the search in a contact-before-motion style [147]. Search-based methods are commonly used for gait planning in legged robot locomotion [148, 149, 150, 151]. Expanding their capabilities in more intricate multi-contact loco-manipulation planning, [152] seeks to implement a graph search method for humanoid grasp contact planning and replanning, and [153] introduces a contact-before-motion planner for multi-contact behaviors.
@@ IV-B Optimization-based Contact Planning
Due to the inherently large problem size, CITOs often rely on speed-up strategies, such as warm-starting with reasonable initial guesses for fast convergence [164] and separating into contact planning and whole-body motion planning subproblems in a hierarchical fashion [165]. With the increasing demand for computation efficiency, CITOs have witnessed a rise in computation speed via sequential quadratic programming (SQP) (e.g.[144]), differential dynamic programming (DDP)(e.g.[166]), and iterative linear quadratic regulator (e.g.[167]). These improvements have even enabled the use of CITO in a Model Predictive Control (a.k.a. CI-MPC) framework for real-time planning on quadruped robots [168, 169] and robotic arms [170, 171]. However, for the humanoid robots, applying CITO to loco-manipulation has yet to be achieved.
Migrating such CITO as real-time CI-MPC to humanoid loco-manipulation presents its own set of challenges, including high-dimensional space of optimization variables, complex/undifferentiable contact dynamics models, proper modeling of interaction dynamics, resolution of initial guess, and tedious tuning. Esteban et al. [172] has made an initial effort to provide reasonable reference trajectories with the Hybrid Linear Inverted Pendulum (HLIP) model to allow natural contact behaviors at both hand and foot locations.
@@ IV-C Learning-based Contact Planning
Learning-based approaches have demonstrated promising potential in planning multi-contact tasks, such as using reinforcement learning to generate velocity commands and contact sequences for quadruped locomotion [173, 174]. These learning-based planners are mostly modular, making it possible to form a hierarchical architecture with model-based planners and controllers at the low level. Compared with traditional optimization-based or heuristics-based approaches, learning-based elements enhance the computation efficiency in multi-contact planning. For example, [175] learns to predict the evolution of centroidal dynamics and contact sequence for dynamic humanoid loco-manipulation in under $0.1$ s, achieving a computation speed boost $300$ times faster compared to traditional optimization-based methods.
@@
In addition, learning-based approaches can assist contact prediction, which allows additional information for contact (re)planning in real time. Precise contacts are often hard to obtain from motion capture data, making it challenging to learn directly from data. To synthesize plausible motion, a naive supervised learning approach often leads to objects moving without any contact or significant penetration between the predicted human body and the objects. [176] introduces contact correction and predicts motions relative to the contacts predicted in human-object interaction tasks. [177] separates the contact prediction and whole-body motion prediction by first predicting the contact positions with a large, dynamic object, which are then used as constraints to synthesize whole-body human motion. These models have the potential to serve as a loco-manipulation planner for humanoid robots. [178] learns to find from the video scenes the affordance, i.e., potential contact points, for household object manipulation tasks. These contacts can be used as heuristics for subsequent motion planning.
IV-D Challenges in Multi-contact Planning
Despite the advancements in multi-contact planning, several key challenges remain. First, computational efficiency remains a fundamental bottleneck. Contact planning inherently involves a combinatorial explosion in possible contact mode sequences, which poses major scalability issues. Optimization-based methods, such as CITO, often suffer from local minima and sensitivity to initialization. Search-based methods, though capable of global exploration, are computationally expensive and unsuitable for online replanning. Learning-based approaches, though computationally efficient, struggle with adaptivity and precision in dynamic contact-rich control. A promising direction lies in combining these paradigms to balance global reasoning, dynamic feasibility, and real-time performance. Second, current approaches rely on simplified point contact models, which fail to capture the complexity of real-world loco-manipulation. In practice, especially with full-body contact and manipulation, robots frequently engage in patch contacts that are complex to model with model-based control and planning frameworks. Extending contact modeling beyond point approximations is essential to better align with real physical interactions. Third, most existing methods work around fixed contact mode sequences, which diverge from the unpredictable and rapidly changing nature of real-world contact interactions. Unexpected contact events, which are often hard to observe due to visual occlusion and lack of whole-body tactile sensing, pose serious challenges for accurate contact estimation, control, and online planning. This raises a fundamental question: is precise contact planning always necessary? To enable more adaptable loco-manipulation behaviors in uncertain environments, future research could explore planning, control, and learning strategies that are robust or even agnostic to small, transient changes in contact.